# Geometry of the moduli space of Hermitian-Einstein connections on manifolds with a dilaton

### v1

@@ 2.2.3 Symmetries and the invariance of the dilaton
Taking the Lie derivative of (1.3) along $K$ and using that $V$ is Killing and leaves both $H$ and $P$ invariant, one finds that
^$\widehat{\nabla}_{i}[V,X]^{\flat}_{j}=0\Longrightarrow\widehat{\nabla}_{i}\mathcal{L}_{V}(D_{j}\Phi)\Longrightarrow D_{i}D_{j}(V\Phi)=0\Longrightarrow D^{2}(V\Phi)=0\Longrightarrow V\Phi=0\leavevmode\nobreak\ .$ (2.40)
@@
$\widehat{R}_{ikmn}+\mathrm{cyclic\leavevmode\nobreak\ in}\leavevmode\nobreak\ \leavevmode\nobreak\ (k,\,m,\,n)=-\widehat{\nabla}_{i}H_{kmn}+\frac{1}{2}dH_{ikmn}\leavevmode\nobreak\ .$ (2.42)
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=\widehat{\nabla}_{i}\theta_{j}\leavevmode\nobreak\ ,$ (2.43)
@@
This proposition holds under weaker assumptions. Consider a compact Riemannian manifold $(M^{d},g,H)$ that satisfies the scale invariance condition (1.3) with metric $g$ and torsion 3-form $H$. Suppose that $K^{\flat}=Z^{\flat}+2d\Phi$ is Killing and leaves both $H$ and $P$ invariant, then it also satisfies (2.20). The proof is along similar lines to the one explained above for Hermitian manifolds.
@@
The proposition above can be further refined. For example if $(M^{2n},g,\omega)$ is a strong KT manifold that satisfies (1.3), then it is a consequence of the results of [35] and [23] that it also satisfies (1.6). In such a case $K$ is $\widehat{\nabla}$-covariantly constant. Therefore, $K$ is both Killing and leaves $H$ invariant. Such refinements of proposition 2.1 will be explored in more detail in section 5. Explicit examples include compact group manifolds with a KT structure and equipped with a bi-invariant metric and torsion 3-form [48, 42].
@@ 3 Symmetries of the moduli spaces
Before we proceed to investigate the symmetries of the moduli space $\mathscr{M}_{\mathrm{HE}}(M^{2n})$ induced from those of the underlying manifold $(M^{2n},g,\omega)$ in the presence of dilaton, it is convenient first to carry out the computations in the Gauduchon gauge. This will complement and provide a proof for the results originally stated in [27]. Later in section 4, we shall introduce a dilaton field on $M^{2n}$ and describe the symmetries of $\mathscr{M}_{\mathrm{HE}}(M^{2n})$ in the presence of a dilaton.
@@ 3.1 Vector fields on the moduli space
$d_{A}a_{X}=d_{A}\iota_{X}F-d_{A}^{2}\eta=(d_{A}\iota_{X}+\iota_{X}d_{A})F-[F,\eta]_{\mathfrak{g}}=\mathcal{L}^{A}_{X}F-[F,\eta]_{\mathfrak{g}}\leavevmode\nobreak\ ,$ (3.2)
@@
Let $(M^{2n},g,\omega)$ be a Hermitian manifold and $F$ be the curvature of an Einstein-Hermitian connection $A$. The vector field $a_{X}=\iota_{X}F-d_{A}\eta$ is tanget to $\mathscr{A}^{*}_{\mathrm{HE}}(M^{2n})$ and horizontal provided that
$\mathcal{O}\eta=-F_{ij}\widehat{\nabla}^{i}X^{j}\leavevmode\nobreak\ .$ (3.6)
@@
$D^{i}_{A}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i}-\mathcal{O}\eta=0\leavevmode\nobreak\ .$ (3.7)
@@
^$\widehat{\nabla}_{A}^{i}F_{ij}+\theta^{i}F_{ij}=0\leavevmode\nobreak\ .$ (3.8)
@@
It is clear from the previous lemma that if $X$ is in addition $\widehat{\nabla}$-covariantly constant, then $\eta=0$ and $\iota_{X}F$ is horizontal. This reproduces the result that has already been established in [27].
@@ 3.2 Holomorphic and Killing vector fields on the moduli space
Let $X$ be holomorphic and Killing vector field on $M^{2n}$ and $F$ be the curvature of an Einstein-Hermitian connection. Consider $a_{X}^{h}=\iota_{X}F-d_{A}\eta$ with $\mathcal{O}\eta=-F_{ij}\widehat{\nabla}^{i}X^{j}$, then
$\Theta(a^{h},a_{X}^{h})=-\Theta(a_{X}^{h},a^{h})=\iota_{X}a^{h}+a^{h}\cdot\eta\leavevmode\nobreak\ ,$ (3.10)
@@
$D_{A}^{i}(\iota_{X}F-d_{A}\eta)_{i}+\theta^{i}(\iota_{X}F-d_{A}\eta)_{i}=0\leavevmode\nobreak\ .$ (3.11)
@@
$a\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=[a^{j},X^{i}F_{ij}]_{\mathfrak{g}}+D_{A}^{j}(X^{i}(d_{A}a)_{ij})+\theta^{j}(d_{A}a)_{ij}X^{i}\leavevmode\nobreak\ .$ (3.12)
@@
$\theta^{j}(d_{A}a)_{ij}X^{i}=X^{i}D_{Ai}(\theta^{j}a_{j})-\theta^{j}D_{Aj}(a_{i}X^{i})\leavevmode\nobreak\ ,$ (3.13)
@@
$a\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=-D_{A}^{2}(a_{i}X^{i})-\theta^{j}D_{Aj}(a_{i}X^{i})+2[a^{j},X^{i}F_{ij}]_{\mathfrak{g}}+X^{i}D_{Ai}(D_{Aj}a^{j}+\theta^{j}a_{j})\leavevmode\nobreak\ .$ (3.17)
@@
$a^{h}\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=-D_{A}^{2}(a^{h}_{i}X^{i})-\theta^{j}D_{Aj}(a^{h}_{i}X^{i})+2[a^{hj},X^{i}F_{ij}]_{\mathfrak{g}}\leavevmode\nobreak\ .$ (3.18)
@@
$-a\cdot(D_{A}^{2}\eta+\theta^{i}D_{Ai}\eta)=-D_{A}^{2}(a\cdot\eta)-\theta^{i}D_{Ai}(a\cdot\eta)-2[a^{i},D_{Ai}\eta]_{\mathfrak{g}}-[D_{A}^{i}a_{i}+\theta^{i}a_{i},\eta]_{\mathfrak{g}}\leavevmode\nobreak\ .$ (3.19)
@@
$-a^{h}\cdot(D_{A}^{2}\eta+\theta^{i}D_{Ai}\eta)=-D_{A}^{2}(a^{h}\cdot\eta)-\theta^{i}D_{Ai}(a^{h}\cdot\eta)-2[a^{hi},D_{Ai}\eta]_{\mathfrak{g}}\leavevmode\nobreak\ .$ (3.20)
@@
$\mathcal{O}(\iota_{X}a^{h}+a^{h}\cdot\eta)=2[a^{hi},X^{j}F_{ji}-D_{Ai}\eta]_{\mathfrak{g}}=2[a^{hi},a^{h}_{{X}i}]_{\mathfrak{g}}\leavevmode\nobreak\ .$ (3.21)
Comparing this to the equation (2.29) that determines $\Theta$, we arrive at (3.10) as the operator $\mathcal{O}$ is invertible. ∎
@@
Let $X$ be a holomorphic and Killing vector field on the Hermitian manifold $(M^{2n},g,\omega)$. Then, $\alpha_{X}$ is a holomorphic and Killing vector field on $\mathscr{M}^{*}_{\mathrm{HE}}$.
@@
$=a_{X}^{h}\cdot(a^{h}_{1},a^{h}_{2})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}-([\kern-2.0pt[a_{X}^{h},a^{h}_{1}]\kern-2.0pt]^{h},a^{h}_{2})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}-(a^{h}_{1},[\kern-2.0pt[a_{X}^{h},a^{h}_{2}]\kern-2.0pt]^{h})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}\leavevmode\nobreak\ ,$ (3.23)
@@
$=\int_{M^{2n}}d^{2n}x\sqrt{g}g^{-1}\mathcal{L}_{X}\langle a_{1}^{h}\,,\,a_{2}^{h}\rangle_{\mathfrak{g}}=0\leavevmode\nobreak\ ,$ (3.33)
@@
and similarly for $(\iota_{X}F-d_{A}\eta)\cdot a_{2}^{h}$. The terms involving $[a_{1}^{h},\eta]_{\mathfrak{g}}$ and $[a_{2}^{h},\eta]_{\mathfrak{g}}$ cancel in (3.33) using that $\langle\cdot,\cdot\rangle_{\mathfrak{g}}$ is bi-invariant. It is clear from (3.33) that $\alpha_{{}_{X}}$ is a Killing vector field on the moduli space ${\mathscr{M}}^{*}_{\mathrm{HE}}$.
@@
$d\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}$ $(\alpha_{1},\alpha_{2},\alpha_{3})=\Big{(}\alpha_{1}\cdot\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}(\alpha_{2},\alpha_{3})-\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}([\kern-2.0pt[\alpha_{1},\alpha_{2}]\kern-2.0pt],\alpha_{3})\Big{)}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(\alpha_{1},\alpha_{2},\alpha_{3})$ (3.36)
$=\Big{(}a^{h}_{1}\cdot\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}(a_{2}^{h},a_{3}^{h})-\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}([\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{h},a_{3}^{h})\Big{)}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.37)
$=\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}([\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{v},a_{3}^{h})+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.38)
$=\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle[\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{v}\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.39)
$=-\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle d_{A}\Theta(a_{1}^{h},a_{2}^{h})\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.40)
@@
$\qquad\qquad+\,\int_{M^{2n}}\omega^{n-1}\wedge\langle\Theta(a_{1}^{h},a_{2}^{h}),d_{A}a^{h}_{3}\rangle_{\mathfrak{g}}\Big{)}+\mathrm{cyclic\leavevmode\nobreak\ in\leavevmode\nobreak\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.42)
[…]
where $[\kern-2.0pt[\cdot,\cdot]\kern-2.0pt]^{v}$ is the vertical component of the commutator and so $[\kern-2.0pt[\cdot,\cdot]\kern-2.0pt]^{v}=-d_{A}\Theta(\cdot,\cdot)$. We have also used that $\omega^{n-1}\wedge d_{A}a_{1}^{h}=0$ as $a_{1}^{h}$ is tangent to ${\mathscr{A}_{\mathrm{HE}}}$ and similarly for $a_{2}^{h}$ and $a_{3}^{h}$.
@@
$=\frac{1}{(n-1)!}\int_{M^{2n}}\omega^{n-1}\wedge\big{(}\langle\mathcal{L}_{X}a_{1}^{h}\wedge a^{h}_{2}\rangle_{\mathfrak{g}}+\langle a_{1}^{h}\wedge\mathcal{L}_{X}a^{h}_{2}\rangle_{\mathfrak{g}}\big{)}=0\leavevmode\nobreak\ ,$ (3.51)
@@
$a_{1}^{h}\cdot(\iota_{X}F-d_{A}\eta)=\iota_{X}d_{A}a_{1}^{h}-d_{A}(a_{1}^{h}\cdot\eta)-[a_{1}^{h},\eta]_{\mathfrak{g}}\leavevmode\nobreak\ ,$ (3.52)
and similarly for $a_{2}^{h}\cdot(\iota_{X}F-d_{A}\eta)$. The terms $[a_{1}^{h},\eta]_{\mathfrak{g}}$ and $[a_{2}^{h},\eta]_{\mathfrak{g}}$ that arise as a result of the computation above do not appear in the calculation described in (3.51) because they cancel as a consequence of the bi-invariance of the inner product $\langle\cdot,\cdot\rangle_{\mathfrak{g}}$. Thus, we have shown that $\alpha_{{}_{X}}$ is also holomorphic and this concludes the proof of the theorem. ∎
@@
$[\kern-2.0pt[\alpha_{X},\alpha_{Y}]\kern-2.0pt]\equiv[\kern-2.0pt[a^{h}_{X},a^{h}_{Y}]\kern-2.0pt]^{h}=-\alpha_{[X,Y]}\leavevmode\nobreak\ .$ (3.53)
@@
Given a holomorphic vector field $X$ on $M^{2n}$, there is a second one given by $Z=-IX$. Moreover, $Z$ is Killing provided that $X$ is Killing and $dX^{\flat}$ is a $(1,1)$-form on $M^{2n}$. Furthermore, if $Z$ is holomorphic and Killing, then $[X,Z]=0$ and so $[\kern-2.0pt[\alpha_{X},\alpha_{Z}]\kern-2.0pt]=0$.
@@
It has been shown in [27] that if $X$ is holomorphic and $\widehat{\nabla}$-covariantly constant – the latter condition implies that $X$ is Killing – and $X^{\flat}\wedge\theta$ is a $(1,1)$-form, then $\alpha_{X}$ is $\widehat{\mathcal{D}}$-covariantly constant, where $\widehat{\mathcal{D}}$ is the compatible with the Hermitian structure connection with skew-symmetric torsion on $\mathscr{M}^{*}_{\mathrm{HE}}$.

### v2

@@ 2.2.3 Symmetries and the invariance of the dilaton
The technique that we use to prove this statement is similar to that employed for the proof of Lemma 2.9 in [49], where $P=0$. Taking the Lie derivative of (1.3) along $K$ and using that $V$ is Killing and leaves both $H$ and $P$ invariant, one finds that
$\widehat{\nabla}_{i}[V,X]^{\flat}_{j}=0\Longrightarrow\widehat{\nabla}_{i}\mathcal{L}_{V}(D_{j}\Phi)\Longrightarrow D_{i}D_{j}(V\Phi)=0\Longrightarrow D^{2}(V\Phi)=0\Longrightarrow V\Phi=0\penalty 10000\ .$ (2.40)
@@
$\widehat{R}_{ikmn}+\mathrm{cyclic\penalty 10000\ in}\penalty 10000\ \penalty 10000\ (k,\,m,\,n)=-\widehat{\nabla}_{i}H_{kmn}+\frac{1}{2}dH_{ikmn}\penalty 10000\ .$ (2.42)
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=\widehat{\nabla}_{i}\theta_{j}\penalty 10000\ ,$ (2.43)
@@
This proposition holds under weaker assumptions, see also [49]. Consider a compact Riemannian manifold $(M^{d},g,H)$ that satisfies the scale invariance condition (1.3) with metric $g$ and torsion 3-form $H$. Suppose that $K^{\flat}=Z^{\flat}+2d\Phi$ is Killing and leaves both $H$ and $P$ invariant, then it also satisfies (2.20). The proof is along similar lines to the one explained above for Hermitian manifolds.
@@
The proposition above can be further refined. For example if $(M^{2n},g,\omega)$ is a strong KT manifold that satisfies (1.3), then it is a consequence of the results of [35, 36] and [23] that it also satisfies (1.6). In such a case $K$ is $^\widehat{\nabla}$-covariantly constant. Therefore, $K$ is both Killing and leaves $H$ invariant. Such refinements of Proposition 2.1 will be explored in more detail in section 5. Explicit examples include compact group manifolds with a KT structure and equipped with a bi-invariant metric and torsion 3-form [50, 43].
@@ 3 Symmetries of the moduli spaces
To investigate the symmetries of the moduli space $\mathscr{M}_{\mathrm{HE}}(M^{2n})$ induced from those on $(M^{2n},g,\omega)$, we shall need the proof of three key Lemmas 3.1, 3.2 and 3.3. Lemma 3.1 holds for any KT manifold $(M^{2n},g,\omega)$. To prove 3.2 and 3.3 depends on the invertibility of the operator $\mathcal{O}$. This holds whenever $M^{2n}$ is compact and the metric $g$ is either Gauduchon, $D^{i}\theta_{i}=0$, or satisfies the condition $D^{i}(e^{-2\Phi}K^{\flat}_{i})=0$ with $K^{\flat}=\theta+2d\Phi$. Thus for the lemmas to hold, it is not necessary to assume that $g$ is Gauduchon.
Before we proceed to investigate the symmetries of the moduli space $\mathscr{M}_{\mathrm{HE}}(M^{2n})$ induced from those of the underlying manifold $(M^{2n},g,\omega)$ in the presence of dilaton, it is convenient first to carry out the computations in the Gauduchon gauge. This will complement and provide a proof for the results originally stated in [27]. Later in section 4, we shall introduce a dilaton field on $M^{2n}$ and describe the symmetries of $\mathscr{M}_{\mathrm{HE}}(M^{2n})$ in the presence of a dilaton. For clarity, we shall mention explicitly whenever $g$ is restricted to satisfy the Gauduchon condition.
@@ 3.1 Vector fields on the moduli space
$d_{A}a_{X}=d_{A}\iota_{X}F-d_{A}^{2}\eta=(d_{A}\iota_{X}+\iota_{X}d_{A})F-[F,\eta]_{\mathfrak{g}}=\mathcal{L}^{A}_{X}F-[F,\eta]_{\mathfrak{g}}\penalty 10000\ ,$ (3.2)
@@
Let $(M^{2n},g,\omega)$ be a Hermitian manifold with a Killing and holomorphic vector field $X$, $\mathcal{O}$ be an invertible operator and $F$ be the curvature of an Einstein-Hermitian connection $A$. The vector field $a_{X}=\iota_{X}F-d_{A}\eta$ is tangent to $\mathscr{A}^{*}_{\mathrm{HE}}(M^{2n})$ and horizontal provided that
$\mathcal{O}\eta=-F_{ij}\widehat{\nabla}^{i}X^{j}\penalty 10000\ .$ (3.6)
@@
$D^{i}_{A}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i}-\mathcal{O}\eta=0\penalty 10000\ .$ (3.7)
@@
$\widehat{\nabla}_{A}^{i}F_{ij}+\theta^{i}F_{ij}=0\penalty 10000\ .$ (3.8)
@@
It is clear from the previous lemma that if $X$ is in addition $^\widehat{\nabla}$-covariantly constant, then $\eta=0$ and $\iota_{X}F$ is horizontal. This reproduces the result that has already been established in [27]. Note also that $\mathcal{O}$ is an invertible operator provided that $M^{2n}$ is compact and $g$ either satisfies the Gauduchon condition or satisfies the condition $D^{i}(e^{-2\Phi}K^{\flat}_{i})=0$ with $K^{\flat}=\theta+2d\Phi$.
@@ 3.2 Holomorphic and Killing vector fields on the moduli space
Let $X$ be holomorphic and Killing vector field on $M^{2n}$, $\mathcal{O}$ be an invertible operator and $F$ be the curvature of an Einstein-Hermitian connection. Consider $a_{X}^{h}=\iota_{X}F-d_{A}\eta$ with $\mathcal{O}\eta=-F_{ij}\widehat{\nabla}^{i}X^{j}$, then
$\Theta(a^{h},a_{X}^{h})=-\Theta(a_{X}^{h},a^{h})=\iota_{X}a^{h}+a^{h}\cdot\eta\penalty 10000\ ,$ (3.10)
@@
$D_{A}^{i}(\iota_{X}F-d_{A}\eta)_{i}+\theta^{i}(\iota_{X}F-d_{A}\eta)_{i}=0\penalty 10000\ .$ (3.11)
@@
$a\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=[a^{j},X^{i}F_{ij}]_{\mathfrak{g}}+D_{A}^{j}(X^{i}(d_{A}a)_{ij})+\theta^{j}(d_{A}a)_{ij}X^{i}\penalty 10000\ .$ (3.12)
@@
$\theta^{j}(d_{A}a)_{ij}X^{i}=X^{i}D_{Ai}(\theta^{j}a_{j})-\theta^{j}D_{Aj}(a_{i}X^{i})\penalty 10000\ ,$ (3.13)
@@
$a\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=-D_{A}^{2}(a_{i}X^{i})-\theta^{j}D_{Aj}(a_{i}X^{i})+2[a^{j},X^{i}F_{ij}]_{\mathfrak{g}}+X^{i}D_{Ai}(D_{Aj}a^{j}+\theta^{j}a_{j})\penalty 10000\ .$ (3.17)
@@
$a^{h}\cdot(D_{A}^{i}\iota_{X}F_{i}+\theta^{i}\iota_{X}F_{i})=-D_{A}^{2}(a^{h}_{i}X^{i})-\theta^{j}D_{Aj}(a^{h}_{i}X^{i})+2[a^{hj},X^{i}F_{ij}]_{\mathfrak{g}}\penalty 10000\ .$ (3.18)
@@
$a\cdot(D_{A}^{2}\eta+\theta^{i}D_{Ai}\eta)=-D_{A}^{2}(a\cdot\eta)-\theta^{i}D_{Ai}(a\cdot\eta)-2[a^{i},D_{Ai}\eta]_{\mathfrak{g}}-[D_{A}^{i}a_{i}+\theta^{i}a_{i},\eta]_{\mathfrak{g}}\penalty 10000\ .$ (3.19)
@@
$a^{h}\cdot(D_{A}^{2}\eta+\theta^{i}D_{Ai}\eta)=-D_{A}^{2}(a^{h}\cdot\eta)-\theta^{i}D_{Ai}(a^{h}\cdot\eta)-2[a^{hi},D_{Ai}\eta]_{\mathfrak{g}}\penalty 10000\ .$ (3.20)
@@
$\mathcal{O}(\iota_{X}a^{h}+a^{h}\cdot\eta)=2[a^{hi},X^{j}F_{ji}-D_{Ai}\eta]_{\mathfrak{g}}=2[a^{hi},a^{h}_{{X}i}]_{\mathfrak{g}}\penalty 10000\ .$ (3.21)
Comparing this to the equation (2.29) that determines $\Theta$, we arrive at (3.10) as the operator $\mathcal{O}$ is invertible. This is always the case provided that $M^{2n}$ is compact and either $g$ is Gauduchon or satisfies $D^{i}(e^{-2\Phi}K^{\flat}_{i})=0$ with $K^{\flat}=\theta+2d\Phi$. ∎
@@
Let $X$ be a holomorphic and Killing vector field on the compact Hermitian manifold $(M^{2n},g,\omega)$ and $g$ be the Gauduchon metric. Then, $\alpha_{X}$ is a holomorphic and Killing vector field on $\mathscr{M}^{*}_{\mathrm{HE}}$.
@@
$=a_{X}^{h}\cdot(a^{h}_{1},a^{h}_{2})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}-([\kern-2.0pt[a_{X}^{h},a^{h}_{1}]\kern-2.0pt]^{h},a^{h}_{2})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}-(a^{h}_{1},[\kern-2.0pt[a_{X}^{h},a^{h}_{2}]\kern-2.0pt]^{h})_{{\mathscr{A}}^{*}_{\mathrm{HE}}}\penalty 10000\ ,$ (3.23)
@@
$=\int_{M^{2n}}d^{2n}x\sqrt{g}g^{-1}\mathcal{L}_{X}\langle a_{1}^{h}\,,\,a_{2}^{h}\rangle_{\mathfrak{g}}=0\penalty 10000\ ,$ (3.33)
@@
and similarly for $(\iota_{X}F-d_{A}\eta)\cdot a_{2}^{h}$. The terms involving $a_{1}^{h},\eta]_{\mathfrak{g}}$ and $a_{2}^{h},\eta]_{\mathfrak{g}}$ cancel in (3.33) using that $\langle\cdot,\cdot\rangle_{\mathfrak{g}}$ is bi-invariant. It is clear from (3.33) that $\alpha_{{}_{X}}$ is a Killing vector field on the moduli space ${\mathscr{M}}^{*}_{\mathrm{HE}}$.
@@
$d\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}$ $(\alpha_{1},\alpha_{2},\alpha_{3})=\Big(\alpha_{1}\cdot\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}(\alpha_{2},\alpha_{3})-\Omega_{\mathcal{M}^{*}_{\mathrm{HE}}}([\kern-2.0pt[\alpha_{1},\alpha_{2}]\kern-2.0pt],\alpha_{3})\Big)+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(\alpha_{1},\alpha_{2},\alpha_{3})$ (3.36)
$=\Big(a^{h}_{1}\cdot\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}(a_{2}^{h},a_{3}^{h})-\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}([\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{h},a_{3}^{h})\Big)+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.37)
$=\Omega_{\mathcal{A}^{*}_{\mathrm{HE}}}([\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{v},a_{3}^{h})+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.38)
$=\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle[\kern-2.0pt[a^{h}_{1},a_{2}^{h}]\kern-2.0pt]^{v}\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.39)
$=-\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle d_{A}\Theta(a_{1}^{h},a_{2}^{h})\wedge a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.40)
@@
$\qquad\qquad+\,\int_{M^{2n}}\omega^{n-1}\wedge\langle\Theta(a_{1}^{h},a_{2}^{h}),d_{A}a^{h}_{3}\rangle_{\mathfrak{g}}\Big)+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.42)
$=\frac{1}{(n-1)!}\,\int_{M^{2n}}d\omega^{n-1}\wedge\langle\Theta(a_{1}^{h},a_{2}^{h}),a^{h}_{3}\rangle_{\mathfrak{g}}+\mathrm{cyclic\penalty 10000\ in\penalty 10000\ }(a^{h}_{1},a^{h}_{2},a^{h}_{3})$ (3.43)
where $\kern-2.0pt[\cdot,\cdot]\kern-2.0pt]^{v}$ is the vertical component of the commutator and so $\kern-2.0pt[\cdot,\cdot]\kern-2.0pt]^{v}=-d_{A}\Theta(\cdot,\cdot)$. We have also used that $\omega^{n-1}\wedge d_{A}a_{1}^{h}=0$ as $a_{1}^{h}$ is tangent to $\mathscr{A}_{\mathrm{HE}}$} and similarly for $a_{2}^{h}$ and $a_{3}^{h}$.
@@
$=\frac{1}{(n-1)!}\int_{M^{2n}}\omega^{n-1}\wedge\big(\langle\mathcal{L}_{X}a_{1}^{h}\wedge a^{h}_{2}\rangle_{\mathfrak{g}}+\langle a_{1}^{h}\wedge\mathcal{L}_{X}a^{h}_{2}\rangle_{\mathfrak{g}}\big)=0\penalty 10000\ ,$ (3.51)
@@
$a_{1}^{h}\cdot(\iota_{X}F-d_{A}\eta)=\iota_{X}d_{A}a_{1}^{h}-d_{A}(a_{1}^{h}\cdot\eta)-[a_{1}^{h},\eta]_{\mathfrak{g}}\penalty 10000\ ,$ (3.52)
and similarly for $a_{2}^{h}\cdot(\iota_{X}F-d_{A}\eta)$. The terms $a_{1}^{h},\eta]_{\mathfrak{g}}$ and $a_{2}^{h},\eta]_{\mathfrak{g}}$ that arise as a result of the computation above do not appear in the calculation described in (3.51) because they cancel as a consequence of the bi-invariance of the inner product $\langle\cdot,\cdot\rangle_{\mathfrak{g}}$. Thus, we have shown that $\alpha_{{}_{X}}$ is also holomorphic and this concludes the proof of the theorem. ∎
@@
$\kern-2.0pt[\alpha_{X},\alpha_{Y}]\kern-2.0pt]\equiv[\kern-2.0pt[a^{h}_{X},a^{h}_{Y}]\kern-2.0pt]^{h}=-\alpha_{[X,Y]}\penalty 10000\ .$ (3.53)
@@
Given a holomorphic vector field $X$ on $M^{2n}$, there is a second one given by Z=IXZ=-IX. Moreover, $Z$ is Killing provided that $X$ is Killing and $dX^{\flat}$ is a $(1,1)$-form on $M^{2n}$. Furthermore, if $Z$ is holomorphic and Killing, then $[X,Z]=0$ and so $\kern-2.0pt[\alpha_{X},\alpha_{Z}]\kern-2.0pt]=0$.
@@
It has been shown in [27] that if $X$ is holomorphic and $^\widehat{\nabla}$-covariantly constant – the latter condition implies that $X$ is Killing – and $X^{\flat}\wedge\theta$ is a $(1,1)$-form, then $\alpha_{X}$ is $\widehat{\mathcal{D}}-$covariantly constant, where $\widehat{\mathcal{D}}$ is the compatible with the Hermitian structure connection with skew-symmetric torsion on $\mathscr{M}^{*}_{\mathrm{HE}}$.